OpenStudy (anonymous):

solvve the following equation: 2x-6-(x-6)=-10

OpenStudy (anonymous):

First we multiply out the negative into the paragraph. 2x-6-x+6=-10 6's cancel out 2x - x = -10 x = -10

OpenStudy (anonymous):

Thanks alot!!

OpenStudy (anonymous):

please simplify: (-2a)(3a)(-5b)(-3b)

OpenStudy (anonymous):

We can group the a's together and the b's together. (-2a)(3a) = -6a^2 (-5b)(-3b) = 15b^2 Now put it back together: -(6a^2)(15b^2)

OpenStudy (anonymous):

so what would the final answer be?

OpenStudy (anonymous):

A lot of teachers prefer different levels/methods of simplifications. You can go ahead and multiply out the numbers, and the final answer will be: -90 a^2 b^2

OpenStudy (anonymous):

Maya needs $240 for the Outer Banks Quick Silver Surf Camp. If she saves $15/week from her allowance and already had $60, how long will it take ehr to save enough?